# Bit Threads: From Entanglement to Geometric Entropies

### v1

@@ Abstract
In this work, we attempt to construct bit threads configurations for various backgrounds using expressions from CPS formalism. We find that when the Ryu-Takayanagi surface is same as horizon, such expressions are sufficient. In other cases, it differs by gradient of a harmonic function. We explore its relation to Wald and differential entropy, and re-express the first law of entanglement entropy in terms of bit threads. Inclusion of quantum effects imposes some constraints on the bulk entanglement via the dominant energy condition. We also apply our method to ascertain bit thread configuration in certain dynamical spacetime.
@@ 1 Introduction
Holographic entanglement entropy has emerged as a useful concept in gauge/gravity duality, providing a crucial bridge between quantum information theory and the geometry of spacetime [1]. It takes us one step closer to understanding various issues regarding the quantum nature of gravity, such as the black hole information problem and many aspects of AdS/CFT correspondence. The Ryu-Takayanagi (RT) formula, equates the entanglement entropy $S_{A}$ of a boundary region $A$ in the dual conformal field theory (CFT) to the area of a minimal codimension-2 surface $m(A)$ anchored on $\partial A$ within a constant time slice of the bulk asymptotically Anti-de Sitter (AdS) spacetime [1, 2]:
@@
Recently, Freedman and Headrick [3] introduced a novel, flow-based method for computing holographic entanglement entropy (HEE) that bypasses the need for prior identification of the RT surface. In this approach, the entanglement entropy is obtained by maximising the flux of a divergenceless vector field over the boundary region. These vector field lines, termed “bit threads”, can be interpreted as representing the quantum correlations (or “Bell pairs”) that connect different parts of the boundary [3]. Bit threads, thus provide a geometric representation of entanglement, offering an alternative perspective to the traditional minimal surface prescription. In this framework, the concept of Bell pairs plays a central role in interpreting the physical meaning of bit threads. Bell pairs, which represent maximally entangled two-level quantum systems, serve as the fundamental units of entanglement. In the bit-thread picture, each thread can be thought of as carrying the information of one Bell pair. In other words, the number of bit threads that connect a given boundary region to its complement reflects the number of Bell pairs shared between these regions, thereby quantifying the entanglement entropy. This interpretation is motivated by the idea that entanglement entropy is a measure of the nonlocal quantum correlations between subsystems. By associating each bit thread with a Bell pair, the flow-based prescription provides an intuitive geometrical representation of these correlations. The maximisation of bit thread flux, which is equivalent to finding the minimal surface in the traditional Ryu-Takayanagi prescription, effectively counts the number of Bell pairs across the boundary. While the identification of each individual thread with a specific Bell pair is a heuristic picture, it offers an alternate way to visualise the abstract notion of entanglement in holographic theories. This perspective not only reinforces the connection between geometry and quantum information but also opens avenues for exploring how modifications to the entanglement structure might reflect in the underlying spacetime geometry. In this flow prescription, the HEE is written as,
@@
Here, $\upsilon^{a}$ is the bit threads vector field and at the RT surface $\upsilon^{a}|_{RT}=\mathrm{C}n^{a}$, where $n^{a}$ is the unit normal of the RT surface and $\mathrm{C}=1/4G_{N}$. The equivalence between this max-flow picture and the minimal-area RT formula is guaranteed by a continuous version of the max-flow min-cut theorem, where the RT surface $m(A)$ emerges as the bottleneck, or minimal cut, whose capacity dictates the maximum possible flux (see Appendix A for a technical review).
@@
This elegant flow-based perspective has turned out to be a useful prescription for a variety of problems. It has been generalised to covariant settings, introduced by Headrick and Hubeny [4], accommodates the dynamics of entanglement in non-static spacetimes by employing a family of divergenceless flows consistent with the extremal surface prescription in the Hubeny-Rangamani-Takayanagi (HRT) framework [5] and beyond that. Additionally, bit threads have been applied to study various information theoretic quantities in holography. They have provided insights into the monogamy of mutual information [6], a fundamental constraint on the sharing of entanglement among multiple parties, as well as the entanglement of purification [7, 8, 9], which quantifies correlations in mixed quantum states. The Riemannian bit threads have also been extended to cases like higher curvature gravity [10], quantum bit threads [11, 12], PEE threads [13], and entanglement threads [14] to name some of the few.
@@
Major thrust of this paper is to explore the relationship between bit threads and the rigorous framework of the covariant phase space (CPS) formalism [15, 16]. Building on prior work that constructed perturbative bit threads using CPS methods [17], we investigate whether the HEE bit thread field can be directly identified with conserved currents derived from CPS. Bit threads, constructed in this way has been referred to as canonical bit threads. This formalism provides a covariant phase space approach, and its integration with bit threads offers a promising avenue for deeper insights into the relationship between geometric and entropic quantities in holography. But there are certain challenges when we directly want to obtain the bit thread vector field from some form derived using the CPS formalism. We find out that while CPS inspired vector fields $v^{a}_{cps}$ are divergenceless, these generally do not satisfy the norm bound and saturation conditions required for bit threads. An extra term $v_{\mathrm{d}}^{a}$ (which does not change the cohomology class of the dual form) needs to be incorporated in many cases, except for certain situations like Rindler-AdS geometries where the RT surface coincides with the horizon. But once we obtain the vector field in Rindler like coordinates, we can use coordinate transformations to find a valid thread configuration in the other coordinate system where we initially needed to add $v^{a}_{\mathrm{d}}$ (like Poincaré) without any prior knowledge of the RT surface.
@@
This connection inspires us to explore whether the flow paradigm can account for other geometric entropy concepts. We argue that the Wald entropy (associated with black hole and cosmological horizons) and differential entropy (associated with spacetimes with a hole) can indeed be naturally expressed as the flux of specific CPS-derived flows linked to the relevant Killing vectors. Furthermore, the CPS method allows us to reformulate the first laws governing variations of these entropies in terms of associated thermodynamic flows, offering a unified geometric perspective for entropy.
@@
Finally, we address the incorporation of quantum effect, discussing the constraint it provides on the bulk entanglement via energy conditions. We also outline the challenges and potential strategies for constructing covariant bit threads in certain dynamical spacetimes, useful for studying processes like thermalisation and black hole evaporation.
@@
The paper is organised as follows. We discuss the construction of a flow from the CPS formalism in section 2. Then we move towards a non-holographic direction to reinterpret the black hole entropy, entropy of cosmological horizons and differential entropy in terms of these threads in section 3. In the section 4, we discuss the first laws of the all the above mentioned entropies in terms of bit threads. The quantum corrections and their relation to the dominant energy condition are discussed in section 5. Section 6 considers dynamical settings. Finally, we conclude with the summary of the work and possible future directions in section 7.
@@ 2 Bit threads and CPS
In general, it is difficult to obtain a bit thread configuration unless the given geometry is very symmetric. There have been some proposals of constructing these vector fields using geodesics [18], unit normals to the RT surfaces, level setting approach [17], or as magnetic flows using a modified Biot-Savart law [19, 20]. Usually, we need to know the metric to do calculations. This motivates us to seek a more fundamental approach, potentially one less reliant on the explicit metric and more connected to the geometry, intrinsic symmetries and conserved structures of the gravitational theory. The CPS formalism, which operates naturally with differential forms and conserved currents, presents a promising avenue [^1] Another approach is to use calibrations, in which the RT surface is realised as a special Lagrangian cycle [21].. So, a natural question arises: Can we construct flows in a more canonical way using the CPS method? [^2] An attempt to construct bit thread configuration using the Noether current form was made in [22]. The authors of [17] have already explored this direction for perturbative bit threads, terming them ‘canonical bit threads’. Here, we investigate how this idea can be extended to non-perturbative cases as well. The main trick is to use parametric variations instead of general perturbations within the CPS framework, which works equally well. Sometimes this is referred to as the ‘solution phase space’ method [23]. For completeness, a short review of the CPS formalism is provided in appendix B. Although there are various versions of the CPS formalism, we will mainly refer to that espoused by Iyer-Wald [15], as all of them lead to the same entropy formula, which is of our primary interest. First, we will sketch a general idea of how to get the bit thread vector field from the CPS, and then we will look into its applicability.
In CPS formalism, the spacetime is considered to be a $d$-dimensional manifold and the phase space is a symplectic manifold. The Hamiltonian then is given by integrating the symplectic current $\boldsymbol{\omega}$ (See (B.3)-(B.8)) over a Cauchy slice. This construction is not only restricted to the Hamiltonian, but is true for any conserved charge associated with a specific set of transformations generated by the symmetry generators. It can be seen if one proceeds as follows. The symplectic current is a closed (d1)(d-1)-form, i.e.,
@@
This equation is called the ‘fundamental identity’ of the CPS formalism. The variation of a conserved charge associated with $\zeta$ is given by,
@@
This is basically the equation (B.8) in terms of the (d2)(d-2)-form $\mathbf{k}$.
@@ 2.1 Codim-2 Current
As mentioned earlier, there are different versions of the CPS formalisms, we will look into the codimension-2 current $\mathbf{k}_{\zeta}$ [^3] Sometimes it is called the ‘surface charge’, but we will stick to the codim-2 current as it is convenient to interpret when we construct a flow using it. for two of them namely the Iyer-Wald formalism and the covariant phase space with boundaries.
@@
with $\xi$ being the diffeomorphism generator. Here $\zeta$ is replaced with $\xi$ to differentiate between a diffeomorphism generator and any general transformation. To obtain an explicit expression for $\mathbf{k}$, the other ingredient is the Noether current, which is defined in terms of the diffeomorphism generator $\xi$, symplectic potential $\mathbf{\Theta}_{\rm{IW}}$ and the Lagrangian $\boldsymbol{\mathrm{L}}$ as given in equation (B.6): $\boldsymbol{\mathrm{J}}_{\xi}\equiv\mathbf{\Theta}_{\rm{IW}}(\mathcal{L}_{\xi}\Phi,\Phi)-\xi.\boldsymbol{\mathrm{L}}$. Then the variation of the Noether current $\boldsymbol{\mathrm{J}}_{\xi}$ yields
$\delta\boldsymbol{\mathrm{J}}_{\xi}=\delta\mathbf{\Theta}_{\rm{IW}}(\mathcal{L}_{\xi}\Phi,\Phi)-\mathcal{L}_{\xi}\mathbf{\Theta}_{\rm{IW}}(\delta\Phi,\Phi)+\mathrm{d}\left(\xi.\mathbf{\Theta}_{\rm{IW}}(\delta\Phi,\Phi)\right),$
where Cartan’s first structure equation has been used. Substituting this in the expression of $\boldsymbol{\omega}_{\rm{IW}}$, one gets
@@
Codim-2 current in CPSB
In the covariant phase space formalism with boundaries [24], one starts with an action of the theory which has an extra boundary Lagrangian (B.9) and follows the similar procedure as stated earlier. In this case, the (pre)-symplectic current is defined as,
$\boldsymbol{\omega}_{\rm{HW}}=\delta\left(\mathbf{\Theta}_{\rm{HW}}-\mathrm{d}\boldsymbol{\mathrm{C}}\right)|_{\tilde{\mathcal{P}}}$ (2.9)
@@
where, $\xi^{\mu}$ is the diffeomorphism generator and $X_{\xi}$ is a vector field defined on the configuration space given by, $X_{\xi}=\int d^{d}x\mathcal{L}_{\xi}\Phi^{i}(x)\frac{\delta}{\delta\Phi^{i}}$. In CPSB, the Noether current is expressed as
@@
and the Noether charge is $\boldsymbol{\mathrm{Q}}_{\xi}$ which satisfies $\boldsymbol{\mathrm{J}}_{\xi}=\mathrm{d}\boldsymbol{\mathrm{Q}}_{\xi}$ as $\mathrm{d}\boldsymbol{\mathrm{J}}_{\xi}=0$. Substituting these in the r.h.s of (2.10), one gets
$X_{\xi}.\boldsymbol{\omega}_{\rm{HW}}=\delta\boldsymbol{\mathrm{J}}_{\xi}+\mathrm{d}\left(\delta_{\xi}\boldsymbol{\mathrm{C}}-\delta\left[X_{\xi}.\boldsymbol{\mathrm{C}}\right]-\xi.\mathbf{\Theta}_{\rm{HW}}\right).$
@@
$\mathrm{d}\left(-X_{\xi}.\boldsymbol{\omega}_{\rm{HW}}\right)=\delta\mathrm{d}\boldsymbol{\mathrm{J}}_{\xi}+\mathrm{d}^{2}\left(\delta_{\xi}\boldsymbol{\mathrm{C}}-\delta\left[X_{\xi}.\boldsymbol{\mathrm{C}}\right]-\xi.\mathbf{\Theta}_{\rm{HW}}\right)=0,$
as $\boldsymbol{\mathrm{J}}_{\xi}$ is locally an exact form and $\mathrm{d}^{2}=0$ by definition. It follows immediately that $X_{\xi}.\boldsymbol{\omega}$’ is a closed form. Locally it can be thought as an exact form and can be expressed as an exterior derivative of some codimension-2 quantity, i.e., $X_{\xi}.\boldsymbol{\omega}_{\rm{HW}}=\mathrm{d}\tilde{\mathbf{k}}_{\xi}$, where $\tilde{\mathbf{k}}_{\xi}$ is given by,
$\tilde{\mathbf{k}}_{\xi}=\delta\boldsymbol{\mathrm{Q}}_{\xi}+\delta_{\xi}\boldsymbol{\mathrm{C}}-\delta\left[X_{\xi}.\boldsymbol{\mathrm{C}}\right]-\xi.\mathbf{\Theta}_{\rm{HW}}\,.$
We can further simplify it by using the boundary conditions and Cartan’s equation, which results in
$\tilde{\mathbf{k}}_{\xi}=\delta\left(\boldsymbol{\mathrm{Q}}_{\xi}+\xi.\boldsymbol{{\ell}}-X_{\xi}.\boldsymbol{\mathrm{C}}\right).$ (2.11)
@@
$\tilde{\boldsymbol{\mathit{j}}_{\xi}}=\left(\boldsymbol{\mathrm{Q}}_{\xi}+\xi.\boldsymbol{{\ell}}-X_{\xi}.\boldsymbol{\mathrm{C}}\right).$ (2.12)
This codimension-2 current can be used to construct various conserved charges. Now we will look how these current forms can be used to construct a flow for bit threads.

### v2

@@ Abstract
In this work, we attempt to construct bit thread configurations for various backgrounds using expressions from the covariant phase space formalism. We find that when the Ryu-Takayanagi surface is same as the horizon, such expressions are sufficient. In other cases, it differs by gradient of a harmonic function. We explore its relation to Wald and differential entropy, and re-express the first law of entanglement entropy in terms of bit threads. Inclusion of quantum effects imposes some constraints on the bulk entanglement via the dominant energy condition. We also apply our method to ascertain a bit thread configuration in a certain dynamical spacetime.
@@ 1 Introduction
Holographic entanglement entropy has emerged as a useful concept in gauge/gravity duality, providing a crucial bridge between quantum information theory and the geometry of spacetime [53]. It takes us one step closer to understanding various issues regarding the quantum nature of gravity, such as the black hole information problem and many aspects of the AdS/CFT correspondence. The Ryu-Takayanagi (RT) formula equates the entanglement entropy $S_{A}$ defined in the dual conformal field theory (CFT) on a boundary region $A$ to the area of a minimal codimension-2 surface $m(A)$ (homologous to $A$) of the bulk asymptotically Anti-de Sitter (AdS) spacetime [53, 52]:
@@
Recently, Freedman and Headrick [27] introduced a novel, flow-based method for computing holographic entanglement entropy (HEE) that bypasses the need for prior identification of the RT surface. In this approach, the entanglement entropy is obtained by maximising the flux of a divergenceless vector field over the boundary region. These vector field lines, termed “bit threads”, can be interpreted as representing the quantum correlations (or “Bell pairs”) that connect different parts of the boundary [27]. Bit threads thus provide a geometric representation of entanglement, offering an alternative perspective to the traditional minimal surface prescription. In this framework, the concept of Bell pairs plays a central role in interpreting the physical meaning of bit threads. Bell pairs, which represent maximally entangled two-level quantum systems, serve as the fundamental units of entanglement. In the bit thread picture, each thread can be thought of as carrying the information of one Bell pair. In other words, the number of bit threads that connect a given boundary region to its complement reflects the number of Bell pairs shared between these regions [^1] It has been shown for pure bipartite states in the context of tensor networks in [45]., thereby quantifying the entanglement entropy. This interpretation is motivated by the idea that entanglement entropy is a measure of the nonlocal quantum correlations between subsystems. By associating each bit thread with a Bell pair, the flow-based prescription provides an intuitive geometrical representation of these correlations. The maximisation of bit thread flux, which is equivalent to finding the minimal surface in the traditional Ryu-Takayanagi prescription, effectively counts the number of Bell pairs across the boundary. While the identification of an individual thread with a specific Bell pair is a heuristic picture, it offers an alternate way to visualise the abstract notion of entanglement in holographic theories. This perspective not only reinforces the connection between geometry and quantum information but also opens avenues for exploring how modifications to the entanglement structure might reflect in the underlying spacetime geometry. In this flow prescription, the HEE is written as,
@@
Here, $\upsilon^{a}$ is the bit thread vector field, which at the RT surface saturates the above inequality: $\upsilon^{a}|_{RT}=\mathrm{C}n^{a}$; $n^{a}$ being the unit normal of the RT surface and $\mathrm{C}=1/4G_{N}$. The equivalence between this max-flow picture and the minimal-area RT formula is guaranteed by a continuous version of the max-flow min-cut theorem, where the RT surface $m(A)$ emerges as the bottleneck, or minimal cut, whose capacity dictates the maximum possible flux (see Appendix A for a technical review).
@@
This elegant, flow-based perspective has turned out to be a useful approach for a variety of problems. It has been generalised to covariant settings, introduced by Headrick and Hubeny [34], which accommodates the dynamics of entanglement in non-static spacetimes by employing a family of divergenceless flows consistent with the extremal surface prescription in the Hubeny-Rangamani-Takayanagi (HRT) framework [37] and beyond. Additionally, bit threads have been applied to study various information-theoretic quantities in holography. They have provided insights into the monogamy of mutual information [20], a fundamental constraint on the sharing of entanglement among multiple parties; as well as the entanglement of purification [23, 8, 32], which quantifies correlations in mixed quantum states. The Riemannian bit threads have also been extended to cases like higher curvature gravity [31], quantum bit threads [4, 51], PEE threads [42], and entanglement threads [44, 43], to name some of the few.
@@
The major thrust of this paper is to explore the relationship between bit threads and the rigorous framework of the covariant phase space (CPS) formalism [39, 58]. Building on prior work that constructed perturbative bit threads using CPS methods [2], we investigate whether the HEE bit thread field can be directly identified with the conserved currents derived from CPS. Bit threads, constructed in this way, have been referred to as canonical bit threads. Since the covariant phase space formalism is very general, bit threads written in this language offer a promising avenue for deeper insights into the geometric interpretation of concepts related to holographic entropy. However, there are certain challenges when we directly attempt to obtain the bit thread vector field from a current form derived using the CPS formalism. We find out that while CPS-inspired vector fields $v^{a}_{cps}$ are divergenceless, these generally do not satisfy the norm bound and saturation conditions required for bit threads. An extra term $v_{\mathrm{d}}^{a}$ (which does not change the cohomology class of the dual form) needs to be incorporated in many cases, except for certain situations like Rindler-AdS geometries, where the RT surface coincides with the horizon. But once we obtain the vector field in Rindler like coordinates, we can use coordinate transformations to find a valid thread configuration in the other coordinate system where we initially needed to add $v^{a}_{\mathrm{d}}$ (like Poincaré) without any prior knowledge of the RT surface.
@@
This connection inspires us to explore whether the flow paradigm can account for other geometric entropy concepts. We argue that the Wald entropy (associated with black holes) and differential entropy (associated with spacetimes with a hole) can indeed be naturally expressed as the flux of specific CPS-derived flows linked to the relevant Killing vectors. Furthermore, the CPS method allows us to reformulate the first laws governing variations of these entropies in terms of associated thermodynamic flows, offering a unified geometric perspective for entropy.
@@
Finally, we address the incorporation of quantum effects, discussing the constraints it provides on bulk entanglement via energy conditions. We also outline the challenges and potential strategies for constructing covariant bit threads in certain dynamical spacetimes, which are helpful in studying processes such as thermalisation and black hole evaporation.
@@
The paper is organised as follows. We discuss the construction of a flow from the CPS formalism in section 2. Then we apply the method to reinterpret the black hole entropy and differential entropy in terms of these threads in certain gravitational backgrounds in section 3. In section 4, we discuss the first laws of the entropies mentioned above in terms of bit threads. The quantum corrections and their relation to the dominant energy condition are addressed in section 5. In section 6, we discuss the bit thread construction in AdS-Vaidya spacetime. Finally, we conclude with the summary of the work and possible future directions in section 7.
@@ 2 Bit threads and CPS
In general, it is difficult to obtain a bit thread configuration unless the given geometry is highly symmetric. There have been proposals for constructing these vector fields using geodesics [3]; unit normals to the RT surfaces, a level-setting approach [2]; or as magnetic flows using a modified Biot-Savart law [50, 28]. Usually, we need to know the metric to do calculations. This motivates us to seek a more fundamental approach, potentially one that is less reliant on the explicit metric and more connected to the geometry, intrinsic symmetries, and conserved structures of the gravitational theory. The CPS formalism, which operates naturally with differential forms and conserved currents, presents a promising avenue. [^2] Another approach is to use calibrations, in which the RT surface is realised as a special Lagrangian cycle [5]. So, a natural question arises: Can we construct flows in a more canonical way using the CPS method? [^3] An attempt to construct a bit thread configuration using the Noether charge form was made in [48]. The authors of [2] have already explored this direction for perturbative bit threads, terming them ‘canonical bit threads’. Here, we investigate how this idea can be extended to non-perturbative cases as well. The main trick is to use parametric variations instead of general perturbations within the CPS framework, which works equally well. Sometimes this is referred to as the ‘solution phase space’ method [29]. For completeness, a short review of the CPS formalism is provided in appendix B. Although there are various versions of the CPS formalism, we will mainly refer to that espoused by Iyer and Wald [39], as all of them lead to the same entropy formula, which is of our primary interest. First, we will sketch the general idea of how to obtain the bit thread vector field from the CPS, and then we will look into its applicability.
In the CPS formalism, the spacetime is considered to be a $d$-dimensional manifold, and the phase space is a symplectic manifold. The Hamiltonian then is given by integrating the symplectic current $\boldsymbol{\omega}$ (See (B.3)-(B.8)) over a Cauchy slice. This construction is not only restricted to the Hamiltonian, but is true for any conserved charge associated with a specific set of transformations generated by the symmetry generators as follows. The symplectic current is a closed (d1;2)(d-1;2)-form [^4] In this notation, a $(p;q)$-form means it is a $p$-form in spacetime and $q$-form in phase space following [29]. We will frequently use the convention $p$-form, which is in the spacetime unless specified., i.e.,
@@
The above equation is called the ‘fundamental identity’ of the CPS formalism [36]. Variation of a conserved charge associated with $\zeta$ is given by,
@@
It is the same equation (B.8) in terms of the (d2)(d-2)-form $\mathbf{k}$.
@@ 2.1 Codim-2 Current
As mentioned earlier, there are different versions of the CPS formalisms, we will look into the codimension-2 current $\mathbf{k}_{\zeta}$, [^5] Sometimes it is called the ‘surface charge’, but we will stick to the terminology of codim-2 current as it is convenient to interpret when we construct a flow using it. for two of them namely, the Iyer-Wald formalism and the covariant phase space with boundaries.
@@
with $\xi$ being the diffeomorphism generator and $\mathcal{L}_{\xi}$ is the Lie derivative w.r.t $\xi$. Here $\zeta$ is replaced with $\xi$ to differentiate between a diffeomorphism generator and any general transformation. To obtain an explicit expression for $\mathbf{k}$, the other ingredient is the Noether current, which is defined in terms of the diffeomorphism generator $\xi$, symplectic potential $\mathbf{\Theta}_{\rm{IW}}$ and the Lagrangian $\boldsymbol{\mathrm{L}}$ as given in equation (B.6): $\boldsymbol{\mathrm{J}}_{\xi}\equiv\mathbf{\Theta}_{\rm{IW}}(\mathcal{L}_{\xi}\Phi,\Phi)-\xi.\boldsymbol{\mathrm{L}}$. Then the variation of the Noether current $\boldsymbol{\mathrm{J}}_{\xi}$ yields
$\delta\boldsymbol{\mathrm{J}}_{\xi}=\delta\mathbf{\Theta}_{\rm{IW}}(\mathcal{L}_{\xi}\Phi,\Phi)-\mathcal{L}_{\xi}\mathbf{\Theta}_{\rm{IW}}(\delta\Phi,\Phi)+\mathrm{d}\left(\xi.\mathbf{\Theta}_{\rm{IW}}(\delta\Phi,\Phi)\right)$
using the Cartan’s magic formula: ${\cal{L}}_{V}\boldsymbol{\alpha}=V.\mathrm{d}\boldsymbol{\alpha}+\mathrm{d}(V.\boldsymbol{\alpha})$ for any vector $V$ and differential-form $\boldsymbol{\alpha}$. Substituting the above expression of $\delta\boldsymbol{\mathrm{J}}_{\xi}$ in the definition of $\boldsymbol{\omega}_{\rm{IW}}$, one gets
@@
Codim-2 current in CPS with boundaries
In the covariant phase space formalism with boundaries [30], one starts with an action of the theory which has an extra boundary Lagrangian (B.9) and follows a similar procedure as stated earlier. In this case, the (pre)-symplectic current is defined as,
$\boldsymbol{\omega}_{\rm{HW}}=\delta\left(\mathbf{\Theta}_{\rm{HW}}-\mathrm{d}\boldsymbol{\mathsf{C}}\right)|_{\tilde{\mathcal{P}}}$ (2.9)
@@
where $\xi^{\mu}$ is the diffeomorphism generator and $X_{\xi}$ is a vector field defined on the configuration space given by, $X_{\xi}=\int d^{d}x\mathcal{L}_{\xi}\Phi^{i}(x)\frac{\delta}{\delta\Phi^{i}}$. In CPSB, the Noether current is expressed as
@@
and the Noether charge is $\boldsymbol{\mathrm{Q}}_{\xi}$ which satisfies $\boldsymbol{\mathrm{J}}_{\xi}=\mathrm{d}\boldsymbol{\mathrm{Q}}_{\xi}$ as $\mathrm{d}\boldsymbol{\mathrm{J}}_{\xi}=0$. Substituting them in the r.h.s of (2.10), one gets
$X_{\xi}.\boldsymbol{\omega}_{\rm{HW}}=\delta\boldsymbol{\mathrm{J}}_{\xi}+\mathrm{d}\left(\delta_{\xi}\boldsymbol{\mathsf{C}}-\delta\left[X_{\xi}.\boldsymbol{\mathsf{C}}\right]-\xi.\mathbf{\Theta}_{\rm{HW}}\right).$
@@
$\mathrm{d}\left(-X_{\xi}.\boldsymbol{\omega}_{\rm{HW}}\right)=\delta\mathrm{d}\boldsymbol{\mathrm{J}}_{\xi}+\mathrm{d}^{2}\left(\delta_{\xi}\boldsymbol{\mathsf{C}}-\delta\left[X_{\xi}.\boldsymbol{\mathsf{C}}\right]-\xi.\mathbf{\Theta}_{\rm{HW}}\right)=0,$
as $\boldsymbol{\mathrm{J}}_{\xi}$ is locally an exact form and $\mathrm{d}^{2}=0$ by definition. It follows immediately that $X_{\xi}.\boldsymbol{\omega}$’ is a closed form. Locally, it can be made an exact form by Poincaré’s lemma, i.e., it can be expressed as an exterior derivative of some codim-2 form: $X_{\xi}.\boldsymbol{\omega}_{\rm{HW}}=\mathrm{d}\tilde{\mathbf{k}}_{\xi}$, where $\tilde{\mathbf{k}}_{\xi}$ is given by,
$\tilde{\mathbf{k}}_{\xi}=\delta\boldsymbol{\mathrm{Q}}_{\xi}+\delta_{\xi}\boldsymbol{\mathsf{C}}-\delta\left[X_{\xi}.\boldsymbol{\mathsf{C}}\right]-\xi.\mathbf{\Theta}_{\rm{HW}}\,.$
We can further simplify it by using the boundary conditions and Cartan’s formula, which results in
$\tilde{\mathbf{k}}_{\xi}=\delta\left(\boldsymbol{\mathrm{Q}}_{\xi}+\xi.\boldsymbol{{\ell}}-X_{\xi}.\boldsymbol{\mathsf{C}}\right).$ (2.11)
@@
$\tilde{\boldsymbol{\mathit{j}}_{\xi}}=\left(\boldsymbol{\mathrm{Q}}_{\xi}+\xi.\boldsymbol{{\ell}}-X_{\xi}.\boldsymbol{\mathsf{C}}\right).$ (2.12)
This codimension-2 current can be used to construct various conserved charges. Now, we will examine how these current forms can be utilised to construct a flow for bit threads.